Acetic acid (HC2H3O2) is an important ingredient of vinegar. A sample of 50.0 mL of a commercial vinegar is titrated against a 1.00 M NaOH solution. What is the concentration (in M) of acetic acid present in the vinegar if 5.75 mL of the base is needed for the titration?
1
Expert's answer
2020-10-05T14:25:33-0400

/CH3COOH + NaOH = CH3COONa + H2O

NCH3COOH * Vsolution CH3COOH = NNaOH * Vsolution NaOH

N = C/feq, feq = 1, N=C.

CCH3COOH * Vsolution CH3COOH = CNaOH * Vsolution NaOH

CCH3COOH = (CNaOH * Vsolution NaOH)/Vsolution CH3COOH

Vsolution NaOH = 5,75 mL = 5,75 * 10-3 L

Vsolution CH3COOH = 50,0 ml = 50,0 * 10-3 L

CCH3COOH = (1,00 * 5,75 *10-3)/(50,0 * 10-3) = 0,115.